# On Generating Monolithic and Model Reconciling Explanations in Probabilistic Scenarios

### v1

@@ 3 A Framework for Probabilistic Explanation Generation
In this section, we outline a framework designed to extend the classical concepts of monolithic explanation, as defined by Definition 1, and model reconciling explanation, as defined by Definition 8, into probabilistic contexts.
@@ 3.1 Probabilistic Monolithic Explanations
Building on the classical notion of monolithic explanation presented in Definition 1, we introduce the concept of a probabilistic monolithic explanation. This concept aims to account for the uncertainty inherent in knowledge bases, providing a framework for explanations that not only identify contributing factors for an explanandum but also quantify the uncertainty in these factors. Throughout this section, we assume a belief base $\mathcal{B}$ and its induced probability distribution $P$.
Formally, a probabilistic monolithic explanation for an explanandum $\varphi$ from belief base $\mathcal{B}$ is defined as follows:
@@
Intuitively, a probabilistic monolithic explanation $\tilde{\epsilon}$ seeks to increase the degree of belief in the explanandum $\varphi$. If $P(\varphi\>|\>\tilde{\epsilon})>P(\varphi)$, this then represents the case where $\tilde{\epsilon}$ increases the degree of belief in $\varphi$ and the greater the value of $P(\varphi\>|\>\tilde{\epsilon})$ the greater the degree of belief in $\varphi$.
@@
Consider the belief base =$\mathcal{B}=\{(a,1),(\lnot a\vee b,2)\}$ and the explanandum $b$. The probability of the explanandum is $P(b)=0.73$. Then, $\tilde{\epsilon}_{1}=\{a\}$ and $\tilde{\epsilon}_{2}=\{\lnot a\vee b\}$ are two probabilistic monolithic explanations for $b$ from $\mathcal{B}$, that is, $P(b\>|\>\tilde{\epsilon}_{1})=0.88>P(b)$ and $P(b\>|\>\tilde{\epsilon}_{2})=0.78>P(b)$.
@@
Looking at Example 4, we can see that monolithic explanations will typically vary in their capacity to increase the degree of belief in the explanandum. In other words, each monolithic explanation provides us with some explanatory gain for the explanandum. Following ? (?, ?), explanatory gain is defined as follows:555 ? (?) originally introduced this measure to quantify the (weak) explanatory power of a hypothesis with respect to evidence, essentially evaluating how effectively the hypothesis explains the evidence.
@@
Now, a natural course of action when seeking monolithic explanations for an explanandum is to seek the one with the highest explanatory gain. While it is tempting to do this, it is important to emphasize that when a monolithic explanation entails the explanandum, then the explanatory gain takes on its greatest value. For example,
@@
Addressing this gap, ? (?) introduced the concept of (strong) explanatory power that integrates the monolithic explanation’s explanatory gain with its probability, offering a more balanced metric for evaluating monolithic explanations. [^8] Good’s measure of (strong) explanatory power is defined as ${{\log}}[\frac{P(\varphi\>|\>h)\cdot P(h)^{\gamma}}{P(\varphi)}]$, where $h$ is a hypothesis and $10<\gamma<$ 1 a constant (?). Building on Good’s measure of explanatory power, we adapt it to our setting and define it as follows: [^9] For a detailed defense of Good’s measure as a quantitative criterion for explanatory power, alongside a discussion of relevant properties and a comprehensive comparison with other measures, we refer the reader to the work by ? (?).
@@
Let $\tilde{\epsilon}$ be a monolithic explanation for explanandum $\varphi$ from belief base $\mathcal{B}$. The explanatory power of $\tilde{\epsilon}$ for $\varphi$ is defined as $\mathrm{EP}(\tilde{\epsilon},\varphi)=G(\tilde{\epsilon},\varphi)+\gamma\cdot P(\tilde{\epsilon})$, where $\gamma\in[0,1]$ is a constant.
@@
Consider the belief base =$\mathcal{B}=\{(a,1.5),(b,3),(\lnot a\vee c,1),(\lnot b\vee c,1)\}$ and the explanandum $c$ with initial probability $P(c)=0.84$. Notice that $\tilde{\epsilon}_{1}=\{a,\lnot a\vee c\}$ and $\tilde{\epsilon}_{2}=\{b,\lnot b\vee c\}$ are two monolithic explanations for $c$ from $\mathcal{B}$, each of which entail $c$ (i.e., $\tilde{\epsilon}_{1}\models c$ and $\tilde{\epsilon}_{2}\models c$), with probabilities $P(\tilde{\epsilon}_{1})=0.68$ and $P(\tilde{\epsilon}_{2})=0.80$, respectively. This means that their explanatory gain for $c$ is equal (Corollary 1), that is, $G(\tilde{\epsilon}_{1},c)=G(\tilde{\epsilon}_{2},c)=0.25$. Now, assuming $\gamma=0.5$, the explanatory power of $\tilde{\epsilon}_{1}$ and $\tilde{\epsilon}_{2}$ respectively is $\mathrm{EP}(\tilde{\epsilon}_{1},c)=0.25+0.5\cdot 0.68=0.59$ and $\mathrm{EP}(\tilde{\epsilon}_{2},c)=0.25+0.5\cdot 0.80=0.65$.
@@
In the next section, we consider how probabilistic monolithic explanations will look like for the model reconciliation problem.
@@ 3.2 Probabilistic Model Reconciling Explanations
Given the knowledge base $\mathrm{KB}_{\alpha}$ of an agent, the belief base $\mathcal{B}_{h}$ of a human user, and an explanandum $\varphi$ such that $\mathrm{KB}_{\alpha}\models\varphi$ and $P_{h}(\varphi)<1$, ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ is a probabilistic model reconciling explanation if and only if $\tilde{\epsilon}^{+}\subseteq\mathrm{KB}_{\alpha}$ and $\tilde{\epsilon}^{-}\subseteq\mathcal{B}_{h}^{\downarrow w}$, and $P_{h}(\varphi\>|\>\tilde{\epsilon}^{+})>P_{h}(\varphi)$ and $P_{h}(\tilde{\epsilon}^{+}\>|\>\lnot\tilde{\epsilon}^{-})>P_{h}(\tilde{\epsilon}^{+})$.
A probabilistic model reconciling explanation ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for $\varphi$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$ is a tuple that increases the degree of belief in $\varphi$ with $\tilde{\epsilon}^{+}$, as well as increasing the degree of belief in $\tilde{\epsilon}^{+}$ with $\tilde{\epsilon}^{-}$ if $\tilde{\epsilon}^{+}$ is inconsistent with $\mathcal{B}_{h}^{\downarrow w}$. For brevity, until the end of this section, we will refer to probabilistic model reconciling explanations $\tilde{\mathcal{E}}$ as model reconciling explanations. In this context, the notion of explanatory gain takes the following form:
@@
Let ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ be a model reconciling explanation for explanandum $\varphi$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$. The explanatory gain of $\tilde{\mathcal{E}}$ for $\varphi$ is defined as $G(\tilde{\mathcal{E}},\varphi)=\log\left(\frac{P(\varphi\>|\>\tilde{\epsilon}^{+})}{P(\varphi)}\right)+\log\left(\frac{P(\tilde{\epsilon}^{+}\>|\>\lnot\tilde{\epsilon}^{-})}{P(\tilde{\epsilon}^{+})}\right)$.
In essence, the explanatory gain of ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for $\varphi$ evaluates to what extent $\tilde{\epsilon}^{+}$ increases the probability of $\varphi$, as well as the extent to which $\tilde{\epsilon}^{-}$ increases the probability of $\tilde{\epsilon}^{+}$, assuming that $\tilde{\epsilon}^{-}$ is false.
@@
Let ~=$\tilde{\mathcal{E}}=\langle\{a,\lnot a\vee b\},\{\lnot c\vee\lnot a\}\rangle$ be the model reconciling explanation for $b$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$ in Example 10. The explanatory gain of $\tilde{\mathcal{E}}$ for $b$ is $G(\tilde{\mathcal{E}},b)=\log\left(\frac{1}{0.5}\right)+\log\left(\frac{0.5}{0.09}\right)=1+2.47=3.47$.
@@
Let ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ be a model reconciling explanation for explanandum $\varphi$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$. The explanatory power of $\tilde{\mathcal{E}}$ for $\varphi$ is defined as $\mathrm{EP}(\tilde{\mathcal{E}},\varphi)=G_{h}(\tilde{\mathcal{E}},\varphi)+\gamma\cdot\left(P_{h}(\tilde{\epsilon}^{+})+P_{h}(\tilde{\epsilon}^{-})\right)$, where $\gamma\in[0,1]$ is a constant.
This definition of explanatory power of ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for $\varphi$ assesses, in addition to the explanatory gain of $\tilde{\mathcal{E}}$, the likelihoods of $\tilde{\epsilon}^{+}$ and $\tilde{\epsilon}^{-}$, with $\gamma$ parameterizing their relative importance in the overall assessment.
@@
Continuing from Example 11, the explanatory power of ~=$\tilde{\mathcal{E}}=\langle\{a,\lnot a\vee b\},\{\lnot c\vee\lnot a\}\rangle$ for $b$ (for $\gamma=0.5$) is $\mathrm{EP}(\tilde{\mathcal{E}},b)=3.47+0.5\cdot(0.09+0.90)=3.96$
@@
Let ~$\tilde{\mathcal{E}}_{1}$ and ~$\tilde{\mathcal{E}}_{2}$ be two model reconciling explanations for explanandum $\varphi$ from knowledge base $\mathrm{KB}_{\alpha}$ for beleif base $\mathcal{B}_{h}$. ~$\tilde{\mathcal{E}}_{1}$ is preferred over ~$\tilde{\mathcal{E}}_{2}$, denoted ~$\tilde{\mathcal{E}}_{1}\succeq\tilde{\mathcal{E}}_{2}$, if and only if $\mathrm{EP}(\tilde{\mathcal{E}}_{1})\geq\mathrm{EP}(\tilde{\mathcal{E}}_{2})$.
@@
Let $\tilde{E}(\varphi)$ denote the set of all model reconciling explanations for explanandum $\varphi$ from knowledge base $\mathrm{KB}_{\alpha}$ for belief base $\mathcal{B}_{h}$. A model reconciling explanation ~$\tilde{\mathcal{E}}^{*}\in\tilde{E}(\varphi)$ is the most-preferred model reconciling explanation for $\varphi$ if and only if $\tilde{\mathcal{E}}^{*}\succeq\tilde{\mathcal{E}}$ for all ~$\tilde{\mathcal{E}}^{\in}\tilde{E}(\varphi)$.

### v2

@@ 5 A Framework for Probabilistic Explanation Generation
In this section, we approach the notion of probabilistic explanation in the following setting, where $\mathcal{B}$ is a belief base and $P$ its induced probability distribution:
Definition 9 (Probabilistic Explanation Generation Problem (PEGP)).
Given a belief base $\mathcal{B}$ and an explanandum $\varphi$ in logic $\mathcal{L}$, the goal of PEGP is to identify an explanation (i.e., a set of formulae) $\tilde{\epsilon}\subseteq\mathcal{B}$ such that the probability of the explanandum given the explanation is higher than the probability of the explanandum, i.e., $P(\varphi\>|\>\tilde{\epsilon})>P(\varphi)$.
[…]
In what follows, we present a framework designed to extend the classical concepts of monolithic explanation, as defined by Definition 1, and model reconciling explanation, as defined by Definition 8, into our probabilistic setting.
@@ 5.1 Probabilistic Monolithic Explanations
Building on the classical notion of monolithic explanation presented in Definition 1, a probabilistic monolithic explanation for an explanandum $\varphi$ from belief base $\mathcal{B}$ is defined as follows:
@@
Consider the belief base $\mathcal{B}=\{(a,1),(\lnot a\vee b,2)\}$ and the explanandum $b$. The probability of the explanandum is $P(b)=0.73$. Then, $\tilde{\epsilon}_{1}=\{a\}$ and $\tilde{\epsilon}_{2}=\{\lnot a\vee b\}$ are two probabilistic monolithic explanations for $b$ from $\mathcal{B}$, that is, $P(b\>|\>\tilde{\epsilon}_{1})=0.88>P(b)$ and $P(b\>|\>\tilde{\epsilon}_{2})=0.78>P(b)$.
@@
Looking at Example 4, we can see that monolithic explanations will typically vary in their capacity to increase the degree of belief in the explanandum. In other words, each monolithic explanation provides us with some explanatory gain for the explanandum. Following ? (?, ?), explanatory gain is defined as follows:888 ? (?) originally introduced this measure to quantify the (weak) explanatory power of a hypothesis with respect to evidence, essentially evaluating how effectively the hypothesis explains the evidence.
@@
A natural course of action when seeking monolithic explanations for an explanandum is to seek the one with the highest explanatory gain. While tempting, it is important to emphasize that when a monolithic explanation entails the explanandum, then the explanatory gain takes on its greatest value. For example,
@@
Addressing this gap, ? (?) introduced the concept of (strong) explanatory power that integrates the monolithic explanation’s explanatory gain with its probability, offering a more balanced metric for evaluating monolithic explanations. [^11] Good’s measure of (strong) explanatory power is defined as ${{\log}}\left(\frac{P(\varphi\>|\>h)\cdot P(h)^{\gamma}}{P(\varphi)}\right)$, where $h$ is a hypothesis and $0<\gamma<1$ a constant (?). Building on Good’s measure of explanatory power, we adapt it to our setting and define it as follows: [^12] For a detailed defense of Good’s measure as a quantitative criterion for explanatory power, alongside a discussion of relevant properties and a comprehensive comparison with other measures, we refer the reader to the work by ? (?).
@@
[…]
@@
Consider the belief base $\mathcal{B}=\{(a,1.5),(b,3),(\lnot a\vee c,1),(\lnot b\vee c,1)\}$ and the explanandum $c$ with initial probability $P(c)=0.84$. Notice that $\tilde{\epsilon}_{1}=\{a,\lnot a\vee c\}$ and $\tilde{\epsilon}_{2}=\{b,\lnot b\vee c\}$ are two monolithic explanations for $c$ from $\mathcal{B}$, each of which entail $c$ (i.e., $\tilde{\epsilon}_{1}\models c$ and $\tilde{\epsilon}_{2}\models c$), with probabilities $P(\tilde{\epsilon}_{1})=0.68$ and $P(\tilde{\epsilon}_{2})=0.80$, respectively. This means that their explanatory gain for $c$ is equal (Corollary 1), that is, $G(\tilde{\epsilon}_{1},c)=G(\tilde{\epsilon}_{2},c)=0.25$. Now, assuming $\gamma=0.5$, the explanatory power of $\tilde{\epsilon}_{1}$ and $\tilde{\epsilon}_{2}$ respectively is $\mathrm{EP}(\tilde{\epsilon}_{1},c)=0.25+0.5\cdot 0.68=0.59$ and $\mathrm{EP}(\tilde{\epsilon}_{2},c)=0.25+0.5\cdot 0.80=0.65$.
@@
5.1.1 Motivating Application: Illustrative Example
Recall from our motivating scenario in Section 4 that we are interested in explaining the explanandum $\varphi=\neg\texttt{move}(room1,room2,A)$, which is akin to asking “why didn’t you take corridor A?” Using our framework, we identify two potential explanations for this explanandum:
  1. 1.
    $\tilde{\epsilon}_{1}=\{\texttt{crowded}(A)\}$ or, in natural language, “Corridor A is very likely crowded.”
    • $P(\varphi|\tilde{\epsilon}_{1})=0.62$
    • Explanatory gain: $G(\tilde{\epsilon}_{1},\varphi)=0.1$
    • Explanatory power: $EP(\tilde{\epsilon}_{1},\varphi)=0.5$
  2. 2.
    $\tilde{\epsilon}_{2}=\{\texttt{crowded}(A),\texttt{move}(room1,room2,A)_{0}\wedge\texttt{crowded}(A)\rightarrow\texttt{robot}\textrm{-}\texttt{at}(room1)_{1},\texttt{move}(room1,room2,A)_{0}\wedge\texttt{crowded}(A)\rightarrow\texttt{robot}\textrm{-}\texttt{at}(room2)_{1}\}$ or, in natural language, “Corridor A is very likely crowded, and when attempting to move through a crowded corridor, my movement outcomes are uncertain; I may either succeed or fail to move through to the next room.”
    • $P(\varphi|\tilde{\epsilon}_{2})=1$
    • Explanatory gain: $G(\tilde{\epsilon}_{2},\varphi)=0.79$
    • Explanatory power: $EP(\tilde{\epsilon}_{2},\varphi)=1.04$
Comparing the two explanations, we see that explanation $\tilde{\epsilon}_{2}$ has the highest explanatory power (1.04), and is thus the most-preferred explanation according to Definition 14. Such an explanation can help the human supervisor to understand both the robot’s belief about the environment (crowded corridor) and how this belief influences action outcomes (movement outcomes under crowded conditions), thereby providing a more complete picture of the robot’s decision-making process.
Comparison with MPE: It is important to highlight the difference with the traditional Most Probable Explanation (MPE) method. MPE typically consists of finding the world with the highest probability given some evidence (?). However, a world does not show the chain of inferences of a given explanandum, and importantly, it is not minimal by definition since it usually includes a (possibly large) number of probabilistic facts whose truth value is irrelevant for the explanandum. For example, applying standard MPE to our running example we get the following outcome:
$\omega_{\text{MPE}}=\{$ $\texttt{crowded}(A),\lnot\texttt{crowded}(B),\lnot\texttt{deliver}_{0},\texttt{deliver}_{1},\lnot\texttt{move}(room1,room2,A)_{0},$
$\lnot\texttt{move}(room1,room2,A)_{1},\texttt{move}(room1,room2,B)_{0},\lnot\texttt{move}(room1,room2,B)_{1},$
$\lnot\texttt{move}(room2,room1,A)_{0},\lnot\texttt{move}(room2,room1,A)_{1},\lnot\texttt{move}(room2,room1,B)_{0},$
$\lnot\texttt{move}(room2,room1,B)_{1},\lnot\texttt{package-delivered}_{0},\lnot\texttt{package-delivered}_{1},$
$\texttt{package-delivered}_{2},\texttt{robot-at}(room1)_{0},\lnot\texttt{robot-at}(room1)_{1},$
$\lnot\texttt{robot-at}(room1)_{2},\lnot\texttt{robot-at}(room2)_{0},\texttt{robot-at}(room2)_{1},$
$\texttt{robot-at}(room2)_{2}\}$
While $\omega_{\text{MPE}}$ is indeed the most probable world ($P(\omega_{\text{MPE}})=0.32$) in which the query ($\neg\texttt{move}(room1,room2,A)$) is true, it does not pinpoint to the (minimal) combination of formulae that justify the query.
@@ 5.2 Probabilistic Model Reconciling Explanations
[…]
A probabilistic model reconciling explanation $\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for $\varphi$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$ is a tuple that increases the degree of belief in $\varphi$ with $\tilde{\epsilon}^{+}$, as well as increasing the degree of belief in $\tilde{\epsilon}^{+}$ with $\tilde{\epsilon}^{-}$ if $\tilde{\epsilon}^{+}$ is inconsistent with $\mathcal{B}_{h}^{\downarrow w}$. For brevity, until the end of this section, we will refer to probabilistic model reconciling explanations $\tilde{\mathcal{E}}$ as model reconciling explanations. In this context, the notion of explanatory gain takes the following form:
@@
[…]
In essence, the explanatory gain of $\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for $\varphi$ evaluates to what extent $\tilde{\epsilon}^{+}$ increases the probability of $\varphi$, as well as the extent to which $\tilde{\epsilon}^{-}$ increases the probability of $\tilde{\epsilon}^{+}$, assuming that $\tilde{\epsilon}^{-}$ is false.
@@
Let $\tilde{\mathcal{E}}=\langle\{a,\lnot a\vee b\},\{\lnot c\vee\lnot a\}\rangle$ be the model reconciling explanation for $b$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$ in Example 10. The explanatory gain of $\tilde{\mathcal{E}}$ for $b$ is $G(\tilde{\mathcal{E}},b)=\log\left(\frac{1}{0.5}\right)+\log\left(\frac{0.5}{0.09}\right)=1+2.47=3.47$.
@@
[…]
This definition of explanatory power of $\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for $\varphi$ assesses, in addition to the explanatory gain of $\tilde{\mathcal{E}}$, the likelihoods of $\tilde{\epsilon}^{+}$ and $\tilde{\epsilon}^{-}$, with $\gamma$ parameterizing their relative importance in the overall assessment.
@@
Continuing from Example 11, the explanatory power of $\tilde{\mathcal{E}}=\langle\{a,\lnot a\vee b\},\{\lnot c\vee\lnot a\}\rangle$ for $b$ (for $\gamma=0.5$) is $\mathrm{EP}(\tilde{\mathcal{E}},b)=3.47+0.5\cdot(0.09+0.90)=3.96$
@@
Let $\tilde{\mathcal{E}}_{1}$ and $\tilde{\mathcal{E}}_{2}$ be two model reconciling explanations for explanandum $\varphi$ from knowledge base $\mathrm{KB}_{\alpha}$ for beleif base $\mathcal{B}_{h}$. $\tilde{\mathcal{E}}_{1}$ is preferred over $\tilde{\mathcal{E}}_{2}$, denoted $\tilde{\mathcal{E}}_{1}\succeq\tilde{\mathcal{E}}_{2}$, if and only if $\mathrm{EP}(\tilde{\mathcal{E}}_{1})\geq\mathrm{EP}(\tilde{\mathcal{E}}_{2})$.
@@
Let $\tilde{E}(\varphi)$ denote the set of all model reconciling explanations for explanandum $\varphi$ from knowledge base $\mathrm{KB}_{\alpha}$ for belief base $\mathcal{B}_{h}$. A model reconciling explanation $\tilde{\mathcal{E}}^{*}\in\tilde{E}(\varphi)$ is the most-preferred model reconciling explanation for $\varphi$ if and only if $\tilde{\mathcal{E}}^{*}\succeq\tilde{\mathcal{E}}$ for all $\tilde{\mathcal{E}}^{\in}\tilde{E}(\varphi)$.
5.2.1 Motivating Application: Illustrative Example
Continuing with our motivating example, we will now examine a case where the human’s belief base differs from the robot’s in a way that leads to divergent expectations about the robot’s decision. Particularly, consider a deterministic version of the robot’s belief base $\mathrm{KB}_{\alpha}$ as described in Section 4, and let us define the human’s belief base $\mathcal{B}_{h}$ with the following key differences:
$(\lnot\texttt{crowded}(A),1)$
$(\texttt{crowded}(B),3)$
The human user believes that corridor A is generally unlikely to be crowded and corridor B is very likely crowded—the opposite of the robot’s beliefs. Additionally, the human believes that the robot can successfully move through corridors A and B, whether crowded or not. Given their belief base $\mathcal{B}_{h}$, the expected optimal plan from the human’s perspective would be $\pi_{h}=\langle\texttt{move}(room1,room2,A)_{0},\texttt{deliver}_{1}\rangle$. When the human observes the robot taking corridor B instead, they would naturally question this decision.
The robot, aware of this discrepancy, can then generate a probabilistic model reconciling explanation $\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ to address $\varphi$:
  1. $\tilde{\epsilon}^{+}=\{\texttt{crowded}(A),\texttt{move}(room1,room2,A)_{0}\wedge\texttt{crowded}(A)\rightarrow\texttt{robot}\textrm{-}\texttt{at}(room1)_{1}\}$
  2. $\tilde{\epsilon}^{-}=\{\lnot\texttt{crowded}(A)\}$
    • $P_{h}(\varphi|\tilde{\epsilon}^{+})=0.7$ (increased from $P_{h}(\varphi)=0.4$)
    • $P_{h}(\tilde{\epsilon}^{+}|\neg\tilde{\epsilon}^{-})=1$ (increased from $P_{h}(\tilde{\epsilon}^{+})=0.15$)
    • Explanatory gain: $G(\tilde{\mathcal{E}},\varphi)=3.04$
    • Explanatory power: $EP(\tilde{\mathcal{E}},\varphi)=3.22$
The explanation essentially communicates: “I didn’t take corridor A because it is very likely crowded, contrary to your belief that it’s not crowded. Additionally, when a corridor is crowded, my movement through it will likely fail, causing me to remain in my original location.” What we see here is that if we focus on both environmental factors (crowded corridor) and action dynamics (movement success), we get an explanation that provides a complete account of the robot’s decision-making process, thereby helping the human gain a better understanding of the robot.